OpenStudy (anonymous):

square root of 8 over 8x^2

OpenStudy (anonymous):

\[\sqrt(8/8x^2)\]

OpenStudy (anonymous):

is this right?

OpenStudy (anonymous):

no \[\sqrt{8}\div 8x ^{2}\]

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

sqrt(8) can be written as sqrt(4)* sqrt(2)

OpenStudy (anonymous):

wait what i just want to know how to simplify it. i have teh answer in the book i just need help with the work. thats not the answer.

OpenStudy (anonymous):

sqrt(4) =2 so 2sqrt(2)/8x^2 sqrt(2)/4x^2

OpenStudy (anonymous):

my book says the answer should be 2sqrt(3)/3

OpenStudy (anonymous):

are you sure you're looking at the right problem? because you're not gonna be able to get rid of the x^2 on the bottom

OpenStudy (anonymous):

oh god your right. SORRY!! is sqrt(8)/ sqrt(6)

OpenStudy (anonymous):

okay for that you do the same thing as above. break the 8 up into 4 * 2 and break the 6 up into 3 * 2 . now you have :\[{\sqrt{4}\sqrt{2}} \over \sqrt{2}\sqrt{3}\] now just divide the sqrt(2) out

OpenStudy (anonymous):

\[{2 \over \sqrt{3}} \times {\sqrt{3} \over \sqrt{3}}\]

OpenStudy (anonymous):

this will get rid of your radical on the bottom and change it to 3

OpenStudy (anonymous):

okay i think i get it. your doing a different way then what im used to but it makes sense

OpenStudy (anonymous):

how do you normally do it?

OpenStudy (anonymous):

heres and example our teacher would break sqrt(8) to the sqrt(2x2x2) and sqrt(6) to sqrt(2x3)